# Proximal Oracles for Optimization and Sampling

### v1

@@ Abstract
We consider convex optimization with non-smooth objective function and log-concave sampling with non-smooth potential (negative log density). In particular, we study two specific settings where the convex objective/potential function is either semi-smooth or in composite form as the finite sum of semi-smooth components. To overcome the challenges caused by non-smoothness, our algorithms employ two powerful proximal frameworks in optimization and sampling: the proximal point framework for optimization and the alternating sampling framework (ASF) that uses Gibbs sampling on an augmented distribution. A key component of both optimization and sampling algorithms is the efficient implementation of the proximal map by the regularized cutting-plane method. We establish the iteration-complexity of the proximal map in both semi-smooth and composite settings. We further propose an adaptive proximal bundle method for non-smooth optimization. The proposed method is universal since it does not need any problem parameters as input. Additionally, we develop a proximal sampling oracle that resembles the proximal map in optimization and establish its complexity using a novel technique (a modified Gaussian integral). Finally, we combine this proximal sampling oracle and ASF to obtain a Markov chain Monte Carlo method with non-asymptotic complexity bounds for sampling in semi-smooth and composite settings.
@@ 1 Introduction
The goal of this paper is to develop efficient proximal algorithms to solve optimization problems (1) as well as to draw samples from potentials (2), where both $f$ in (1) and (2) lack smoothness. In particular, we consider two settings where the convex objective/potential function $f$ is either semi-smooth (i.e., the (sub)gradient $f^{\prime}$ is Hölder-continuous with exponent $\alpha\in[0,1]$) or a composite function with multiple semi-smooth components. The core of both proximal optimization and sampling algorithms is a proximal map of $f$. Based on an efficient implementation of the proximal map, we develop an adaptive proximal bundle method to solve (1). Building upon the proximal map of $f$ and rejection sampling, we propose a highly efficient method to realize a proximal sampling oracle, which is used in a proximal sampling framework [23, 5] in the same spirit as the proximal point method for optimization. With those proximal oracles for optimization and sampling in hand, we are finally able to establish the complexity to sample from densities with non-smooth potentials.
@@
    We analyze the complexity bounds for implementing the proximal map of $f$ using the regularized cutting-plane method in both semi-smooth and composite settings (Section 3).
@@
    We develop an adaptive proximal bundle method (APBM) using the regularized cutting-plane method and a novel adaptive stepsize strategy in the proximal point method, and establish the complexity bound for semi-smooth optimization (Section 4). APBM is a universal method as it does not need any problem-dependent parameters as input.
@@
    We propose an efficient scheme to realize the proximal sampling oracle that lacks smoothness and establish novel techniques to bound its complexity. Combining the proximal sampling oracle and the proximal sampling framework, we obtain a general proximal sampling algorithm for convex semi-smooth and composite potentials. Finally, we establish complexity bounds for the proximal sampling algorithm in both cases (Section 5).
@@ 2 Proximal optimization and sampling
Many classical first-order methods in optimization, such as the proximal gradient method, the proximal subgradient method, the primal-dual hybrid gradient method of [4] (also known as the Chambolle-Pock method), the extragradient method of [21] are instances of PPF. It is worth noting that, by showing that the alternating direction method of multipliers (ADMM) as an instance of PPF, [32] gives the first iteration-complexity result of ADMM for solving a class of linearly constrained convex programming problems. Inspired by the PPF viewpoint, [27] develops a variant of the proximal bundle method and establishes the optimal iteration-complexity, which is the first optimal complexity result for proximal bundle methods. Recent works [19, 26, 20, 29] have also applied PPF to solve weakly convex optimization and weakly convex-concave min-max problems.
@@
A rejection sampling-based implementation of RGO for general convex nonsmooth potential function $f$ with bounded Lipschitz constant is given in [24]. If the stepsize $\eta$ is small enough, then it only takes a constant number of rejection steps to generate a sample according to RGO in expectation. Another exact realization of RGO is provided in [25] for nonconvex composite potential $f$ satisfying Hölder continuous conditions. It is also shown that the expected number of rejections to implement RGO is a small constant if $\eta$ is small enough. Other inexact realizations of RGO based on approximate rejection sampling are studied in [14, 10]. In all these implementations, a key step is realizing the proximal map (4). It is worth noting that [25] also connects ASF with other well-known Langevin-type sampling algorithms such as Langevin Monte Carlo (LMC) and Proximal Langevin Monte Carlo (PLMC) via RGO. In a nutshell, [25] shows that both LMC and PLC are instances of ASF but with approximate implementations of RGO, which always accept the sample from the proposal distribution without rejection. Hence, this explains why the samples generated by LMC is biased while ASF is unbiased.
@@
For both optimization and sampling, we consider two specific scenarios: 1) $f$ is semi-smooth, i.e., $f$ satisfies
@@
where $f^{\prime}$ denotes a subgradient of $f$, $\alpha\in[0,1]$, and $L_{\alpha}>0$; and 2) $f$ is a composite function of semi-smooth components, i.e., $f$ satisfies
@@ 3 Algorithm and complexities for the proximal subproblem
The proximal subproblem (3) generally does not admit a closed-form solution. We design an iterative method that approximately solves (3) and derive the corresponding iteration-complexities for semi-smooth and composite $f$ in Subsections 3.1 and 3.2, respectively.
@@
and aim at obtaining a $\delta$-solution, i.e., a point $\bar{x}$ such that $f_{y}^{\eta}(\bar{x})-f_{y}^{\eta}(x^{*})\leq\delta$. In both semi-smooth and composite settings, we use a regularized cutting-plane method (Algorithm 2), which is usually used in the proximal bundle method [27, 28] for solving convex non-smooth optimization problems. We remark that though Algorithm 2 is widely used in the proximal bundle method and is not new, the complexity analyses (i.e., Theorems 3.5 and 3.9) for semi-smooth and composite functions $f$ are lacking.
@@
  1. Let $y\in\mathbb{R}^{d}$, $\eta>0$, and $\delta>0$ be given, and set $x_{0}=\tilde{x}_{0}=y$, and $j=1$.
@@ 3. Compute
  4. If $f_{y}^{\eta}(\tilde{x}_{j})-f_{j}^{\eta}(x_{j})\leq\delta$, then stop and return $J=j,x_{J},\tilde{x}_{J}$; else, go to step 5.
  5. Set $j$leftarrow $j+1$ and go to step 2.
@@
Proof: a) The first inequality follows from the definition of $f_{j}$ in step 2 of Algorithm 2. The second inequality directly follows from the definition of $f_{j}$ and the convexity of $f$.
@@
$\delta_{j+1}$ $=f_{y}^{\eta}(\tilde{x}_{j+1})-f_{j+1}^{\eta}(x_{j+1})\leq f_{y}^{\eta}(\tilde{x}_{j})-f_{j}^{\eta}(x_{j})-\frac{1}{2\eta}\|x_{j+1}-x_{j}\|^{2}=\delta_{j}-\frac{1}{2\eta}\|x_{j+1}-x_{j}\|^{2}.$
@@
e) This statement directly follows from the optimality condition of (9).
@@
3.1 Complexity for semi-smooth optimization
This subsection is devoted to the complexity analysis of Algorithm 2 for solving (9) where $f$ is semi-smooth, i.e., satisfying (5). The following lemma provides basic recursive formulas and is the starting point of the analysis of Algorithm 2.
@@ 3.1 Complexity for semi-smooth optimization
Assume $f$ is convex and $L_{\alpha}$-semi-smooth. Then, for every $j\geq 1$, the following statements hold:
@@
Proof: a) It follows from the definition of $\delta_{j}$ in (12) and the definition of $\tilde{x}_{j}$ in (11) that
@@
b) This statement directly follows from a) and the second inequality in Lemma 3.1(c).
@@
Proof: a) Using the definition of $\beta$ in (13), the assumption that $\delta_{j}>\delta$, and Lemma 3.2(b), we obtain
@@
where the last inequality is due to the definition of $j_{0}$ in (13). This contradicts with the assumption that $\delta_{j_{0}}>\delta$, and hence we prove this statement.
@@
Proof: Following the optimality condition of (10) with $j=1$, we have $x_{0}-x_{1}=\eta f^{\prime}(x_{0})=\eta f^{\prime}(y)$. This identity and Lemma 3.2(a) with $j=1$ then imply that the lemma holds.
@@
Proof: This theorem follows directly from Proposition 3.3 and Lemma 3.4.
3.2 Complexity for composite optimization
This subsection is devoted to the complexity analysis of Algorithm 2 for solving (9) where $f$ is a composite function satisfying (6). The following lemma is an analogue of Lemma 3.2 and provides key recursive formulas for $\delta_{j}$, which is defined in (12).
@@ 3.2 Complexity for composite optimization
Proof: a) Following a similar argument as in the proof of Lemma 3.2(a) with (7) replaced by (8), we have
@@
and hence the statement follows.
@@
Proof: This lemma follows from a similar argument as in the proof of Lemma 3.4.
@@
Hence, it is easy to see that $\delta_{j}\leq\delta$ if $j\geq\frac{1}{1-\tau}\log\left(\frac{2\delta_{1}}{\delta}\right)$. Using the definition of $\tau$ in (17), we have if $j$ is as in (16), then $\delta_{j}\leq\delta$.
@@
Proof: This theorem follows directly from Proposition 3.8 and Lemma 3.7.
@@ 4 Adaptive proximal bundle method
For both the subgradient method and the proximal bundle method to have the optimal performance, one needs to carefully select the stepsize $\eta$, namely, being small enough for the subgradient method and within a certain (but relatively large) range for the proximal bundle method. In both cases, we need to know the problem-dependent parameters such as $\alpha$ and $L_{\alpha}$, which are unknown or hard to estimate in practice. In this section, we develop the APBM based on an adaptive stepsize strategy for (3) and using Algorithm 2 to solve each subproblem (3). We also discuss variants of adaptive subgradient methods and compare them with APBM. For simplicity, we only present the analysis of semi-smooth functions satisfying (5), while the composite functions satisfying (6) can be similarly analyzed using results from Subsection 3.2.
From practical observations, the proximal bundle method works well when the number of inner iterations (i.e., those of Algorithm 2 to solve (3)) stays as a constant much larger than 1 (i.e., that of the subgradient method), say $10$. Recall from Theorem 3.5 that inner complexity is $\tilde{\cal O}\left(\eta L_{\alpha}^{\frac{2}{\alpha+1}}\left(\frac{1}{\delta}\right)^{\frac{1-\alpha}{\alpha+1}}+1\right)$. Since we do not know $\alpha$ and $L_{\alpha}$, we cannot choose a constant stepsize $\eta$ so that the number of inner iterations is close to a desired number such as $10$. Hence, an adaptive stepsize rule is indeed needed.
@@
  1. Let $y_{0}\in\mathbb{R}^{d}$, $\eta_{0}>0$, $\beta_{0}\in(0,1]$, and $\varepsilon>0$ be given, and set $k=1$.
  2. Call Algorithm 2 with $(y,\eta,\delta)=(y_{k-1},\eta_{k-1},\varepsilon/2)$ and output $(y_{k},\tilde{y}_{k})=(x_{J},\tilde{x}_{J})$.
  3. In the execution of Algorithm 2, if (18) is always true, then set $\eta_{k}=\eta_{k-1}$; otherwise, set $\eta_{k}=\eta_{k-1}/2$.
  4. Set $k$leftarrow $k+1$ and go to step 2.
@@
Assume $f$ is convex and $L_{\alpha}$-semi-smooth. The following statements hold for APBM:
@@
Proof: a) It follows from Lemma 3.1(d) that for every $u\in\mathbb{R}^{d}$
@@
d) This statement immediately follows from (b) and step 3 of Algorithm 3.
@@
Assume $f$ is convex and $L_{\alpha}$-semi-smooth. If $\eta_{0}\leq\|y_{0}-x_{*}\|^{2}/\varepsilon$, then the iteration-complexity to obtain an $\varepsilon$-solution to (1) is given by
@@
Proof: Noting from Lemma 4.1 (d) that $\eta_{k}$ is bounded from below, we know there exists some $\tilde{\eta}\in[\underline{\eta},\eta_{0}]$ such that for some $k_{0}\geq 1$, $\eta_{k}\equiv\tilde{\eta}$ for $k\geq k_{0}$. Thus, it follows from the assumption that $\eta_{0}\leq\|y_{0}-x_{*}\|^{2}/\varepsilon$ that
@@
$2\sum_{k=1}^{n}\eta_{k-1}\left(\min_{1\leq k\leq n}f(\tilde{y}_{k})-f(u)\right)\leq 2\sum_{k=1}^{n}\eta_{k-1}[f(\tilde{y}_{k})-f(u)]\leq\|y_{0}-u\|^{2}-\|y_{n}-u\|^{2}+\varepsilon\sum_{k=1}^{n}\eta_{k-1}.$
@@
for cycles $k\leq k_{0}-1$. Finally, the total iteration-complexity (22) clearly follows from the above two bounds.
Discussion on other universal methods Several universal methods based on the backtracking line-search procedure have been studied in the literature. Paper [37] considers the same semi-smooth problem (with an additional composite function $h$) as in this paper. To finds an $\varepsilon$-solution of (1), the universal primal gradient method proposed in [37] starts from an initial pair $(\hat{x}_{0},\eta_{0})$ and in the $(j+1)$-th iteration searches for a pair $(x_{\eta},\eta)$ satisfying a condition
@@
If the condition (29) is not satisfied, then the method rejects the pair, sets $\eta\leftarrow\eta/2$, and updates $x_{\eta}$ as in (30) with the new $\eta$, otherwise, it accepts the pair and sets $(\hat{x}_{j+1},\eta_{j+1})=(x_{\eta},\eta)$. Paper [28] proposes an adaptive composite subgradient (A-CS) method for solving (1) where $f$ satisfies
@@
for any $\theta>0$. Hence, the semi-smooth functions (5) considered in this paper are included in the class of functions satisfying (31). More interestingly, a careful look at A-CS of [28] and the universal primal gradient method of [37] reveals that the two methods are identical.
@@
The well-known Nesterov’s accelerated gradient method has been shown in [36] to match the above complexity bound and hence is an optimal method. On the other hand, the dominant term of bound (22) is its first term and it is only optimal when $\alpha=0$, i.e., $f$ is $L_{0}$-Lipschitz continuous. Motivated by [36], it is possible to develop optimal universal methods based on the accelerated gradient method. This requires accelerated schemes in both PPF and Algorithm 2. Paper [31] proposes an accelerated variant of PPF, which is extended by [3, 16, 11] to obtain optimal $p$-th order methods with convergence rate ${\cal O}(k^{-(3p+1)/2})$ for $p\geq 2$.

### v2

@@ Abstract
We consider convex optimization with non-smooth objective function and log-concave sampling with non-smooth potential (negative log density). In particular, we study two specific settings where the convex objective/potential function is either Hölder smooth or in hybrid form as the finite sum of Hölder smooth components. To overcome the challenges caused by non-smoothness, our algorithms employ two powerful proximal frameworks in optimization and sampling: the proximal point framework for optimization and the alternating sampling framework (ASF) that uses Gibbs sampling on an augmented distribution. A key component of both optimization and sampling algorithms is the efficient implementation of the proximal map by the regularized cutting-plane method. We establish its iteration-complexity under both Hölder smoothness and hybrid settings using novel convergence analysis, yielding results that are new to the literature. We further propose an adaptive proximal bundle method for non-smooth optimization that employs an aggressive adaptive stepsize strategy, which adjusts stepsizes only when necessary and never rejects iterates. The proposed method is universal since it does not need any problem parameters as input. Additionally, we provide an exact implementation of a proximal sampling oracle, analogous to the proximal map in optimization, along with simple complexity analyses for both the Hölder smooth and hybrid cases, using a novel technique based on a modified Gaussian integral. Finally, we combine this proximal sampling oracle and ASF to obtain a Markov chain Monte Carlo method with non-asymptotic complexity bounds for sampling in Hölder smooth and hybrid settings.
@@ 1 Introduction
The goal of this paper is to develop efficient proximal algorithms to solve optimization problems (1) as well as to draw samples from potentials (2), where both $f$ in (1) and (2) lack smoothness. In particular, we consider two settings where the convex objective/potential function $f$ is either Hölder smooth (i.e., the (sub)gradient $f^{\prime}$ is Hölder-continuous with exponent $\alpha\in[0,1]$) or a hybrid function with multiple Hölder smooth components. The core of both proximal optimization and sampling algorithms is a proximal map of $f$. Based on an efficient implementation of the proximal map, we develop an adaptive proximal bundle method to solve (1). Building upon the proximal map of $f$ and rejection sampling, we propose a highly efficient method to realize a proximal sampling oracle, which is used in a proximal sampling framework [29, 5] in the same spirit as the proximal point method for optimization. With those proximal oracles for optimization and sampling in hand, we are finally able to establish the complexity to sample from densities with non-smooth potentials.
@@
    We analyze the complexity bounds for implementing the proximal map of $f$ using the regularized cutting-plane method in both Hölder smooth and hybrid settings (Section 3). The complexity analyses for both Hölder smooth and hybrid cases, presented in Subsections 3.1 and 3.2, respectively, are novel contributions to the literature and employ proof techniques distinct from existing works such as [7, 8, 23, 34, 38].
@@
    We develop an adaptive proximal bundle method (APBM) using the regularized cutting-plane method and a novel adaptive stepsize strategy in the proximal point method, and establish the complexity bound for Hölder smooth optimization (Section 4). APBM is a universal method as it does not need any problem-dependent parameters as input. In contrast to standard universal methods based on conservative line searches on stepsizes, such as the universal primal gradient method of [48], APBM has the benefit of adjusting stepsizes only when necessary and never rejects iterates.
@@
    We propose an efficient scheme to realize the proximal sampling oracle that lacks smoothness and establish novel techniques to bound its complexity. Combining the proximal sampling oracle and the proximal sampling framework, we obtain a general proximal sampling algorithm for convex Hölder smooth and hybrid potentials. Finally, we establish complexity bounds for the proximal sampling algorithm in both cases (Section 5). The complexity bounds presented in Section 5 are similar to those in [12]; however, they are derived under the assumption of an exact proximal sampling oracle, whereas [12] considers an inexact implementation of the oracle. The contributions of Section 5 lie in providing much simpler complexity analyses for the exact realization of the proximal sampling oracle in both the Hölder smooth and hybrid cases, compared to the existing analyses in [35, 36].
It is worth noting that this paper does not aim to establish the optimal complexity of universal methods or to improve the complexity of proximal sampling algorithms. Instead, it develops a regularized cutting-plane method as an efficient implementation of the proximal oracle used in both proximal optimization and sampling, and demonstrates its interesting applications in universal methods and proximal sampling algorithms.
@@ 2 Proximal optimization and sampling
Many classical first-order methods in optimization, such as the proximal gradient method, the proximal subgradient method, the primal-dual hybrid gradient method of [4] (also known as the Chambolle-Pock method), the extra gradient method of [26] are instances of PPF. It is worth noting that, by showing that the alternating direction method of multipliers (ADMM) as an instance of PPF, [43] gives the first iteration-complexity result of ADMM for solving a class of linearly constrained convex programming problems.
Another example of PPF is the proximal bundle method, which was first proposed in [31, 32, 41, 63] and further developed in [7, 8, 13, 23, 34, 38, 49, 56, 60]. Notably, inspired by the PPF viewpoint, papers [34, 38] develop a variant of the proximal bundle method and establish the optimal iteration-complexity, which is the first optimal complexity result for proximal bundle methods. Recent works [24, 37, 25, 39] have also applied PPF to solve weakly convex optimization and weakly convex-concave min-max problems.
@@
A rejection sampling-based implementation of RGO for general convex nonsmooth potential function $f$ with bounded Lipschitz constant is given in [35]. If the stepsize $\eta$ is small enough, then it only takes a constant number of rejection steps to generate a sample according to RGO in expectation. Another exact realization of RGO is provided in [36] for nonconvex hybrid potential $f$ satisfying Hölder continuous conditions. It is also shown that the expected number of rejections to implement RGO is a small constant if $\eta$ is small enough. Other inexact realizations of RGO based on approximate rejection sampling are studied in [17, 12]. In all these implementations, a key step is realizing the proximal map (4). It is worth noting that [36] also connects ASF with other well-known Langevin-type sampling algorithms such as Langevin Monte Carlo (LMC) and Proximal Langevin Monte Carlo (PLMC) via RGO. In a nutshell, [36] shows that both LMC and PLC are instances of ASF but with approximate implementations of RGO, which always accept the sample from the proposal distribution without rejection. Hence, this explains why the samples generated by LMC is biased while ASF is unbiased.
@@
For both optimization and sampling, we consider two specific scenarios: 1) $f$ is Hölder smooth, i.e., $f$ satisfies
@@
where $f^{\prime}$ denotes a subgradient of $f$, $\alpha\in[0,1]$, and $L_{\alpha}>0$; and 2) $f$ is a hybrid function of Hölder smooth components, i.e., $f$ satisfies
@@ 3 Algorithm and complexities for the proximal subproblem
The proximal subproblem (3) generally does not admit a closed-form solution. We design an iterative method that approximately solves (3) and derive the corresponding iteration-complexities for Hölder smooth and hybrid $f$ in Subsections 3.1 and 3.2, respectively.
@@
and aim at obtaining a $\delta$-solution, i.e., a point $\bar{x}$ such that $f_{y}^{\eta}(\bar{x})-f_{y}^{\eta}(x^{*})\leq\delta$. In both Hölder smooth and hybrid settings, we use a regularized cutting-plane method (Algorithm 2), which is usually used in the proximal bundle method [38, 34] for solving convex non-smooth optimization problems. We remark that though Algorithm 2 is widely used in the proximal bundle method and is not new, the complexity analyses (i.e., Theorems 3.5 and 3.9) for Hölder smooth and hybrid functions $f$ are lacking.
@@
0:  Let $y\in\mathbb{R}^{d}$, $\eta>0$, and $\delta>0$ be given, and set $x_{0}=\tilde{x}_{0}=y$.
  for $j=1,2,\cdots$ do
@@
     If $f_{y}^{\eta}(\tilde{x}_{j})-f_{j}^{\eta}(x_{j})\leq\delta$, then stop and return $J=j,x_{J},\tilde{x}_{J}$.
  end for
@@
Proof a) The first inequality follows from the definition of $f_{j}$ in step 2 of Algorithm 2. The second inequality directly follows from the definition of $f_{j}$ and the convexity of $f$.
@@
$\delta_{j+1}$ $=f_{y}^{\eta}(\tilde{x}_{j+1})-f_{j+1}^{\eta}(x_{j+1})\leq f_{y}^{\eta}(\tilde{x}_{j})-f_{j}^{\eta}(x_{j})-\frac{1}{2\eta}\|x_{j+1}-x_{j}\|^{2}$
$=\delta_{j}-\frac{1}{2\eta}\|x_{j+1}-x_{j}\|^{2}.$
@@
e) This statement directly follows from the optimality condition of (9). ∎
@@
3.1 Complexity for Hölder smooth optimization
This subsection is devoted to the complexity analysis of Algorithm 2 for solving (9) where $f$ is Hölder smooth, i.e., satisfying (5). The following lemma provides basic recursive formulas and is the starting point of the analysis of Algorithm 2.
@@ 3.1 Complexity for Hölder smooth optimization
Assume $f$ is convex and $L_{\alpha}$-Hölder smooth. Then, for every $j\geq 1$, the following statements hold:
@@
Proof a) It follows from the definition of $\delta_{j}$ in (13) and the definition of $\tilde{x}_{j}$ in (12) that
@@
b) This statement directly follows from a) and the second inequality in Lemma 3.1(c). ∎
@@
Proof a) Using the definition of $\beta$ in (14), the assumption that $\delta_{j}>\delta$, and Lemma 3.2(b), we obtain
@@
where the last inequality is due to the definition of $j_{0}$ in (14). This contradicts with the assumption that $\delta_{j_{0}}>\delta$, and hence we prove this statement. ∎
@@
Proof Following the optimality condition of (11) with $j=1$, we have $x_{0}-x_{1}=\eta f^{\prime}(x_{0})=\eta f^{\prime}(y)$. This identity and Lemma 3.2(a) with $j=1$ then imply that the lemma holds. ∎
@@
Proof This theorem follows directly from Proposition 3.3 and Lemma 3.4. ∎
3.2 Complexity for hybrid optimization
This subsection is devoted to the complexity analysis of Algorithm 2 for solving (9) where $f$ is a hybrid function satisfying (6). The following lemma is an analogue of Lemma 3.2 and provides key recursive formulas for $\delta_{j}$, which is defined in (13).
@@ 3.2 Complexity for hybrid optimization
Proof a) Following a similar argument as in the proof of Lemma 3.2(a) with (7) replaced by (8), we have
@@
and hence the statement follows. ∎
@@
Proof This lemma follows from a similar argument as in the proof of Lemma 3.4. ∎
@@
Hence, it is easy to see that $\delta_{j}\leq\delta$ if $j\geq\frac{1}{1-\tau}\log\left(\frac{2\delta_{1}}{\delta}\right)$. Using the definition of $\tau$ in (18), we have if $j$ is as in (17), then $\delta_{j}\leq\delta$. ∎
@@
Proof This theorem follows directly from Proposition 3.8 and Lemma 3.7. ∎
@@ 4 Adaptive proximal bundle method
For both the subgradient method and the proximal bundle method to have the optimal performance, one needs to carefully select the stepsize $\eta$, namely, being small enough for the subgradient method and within a certain (but relatively large) range for the proximal bundle method. In both cases, we need to know the problem-dependent parameters such as $\alpha$ and $L_{\alpha}$, which are unknown or hard to estimate in practice. In this section, we develop the APBM based on an adaptive stepsize strategy for (3) and using Algorithm 2 to solve each subproblem (3). We also discuss variants of adaptive subgradient methods and compare them with APBM. For simplicity, we only present the analysis of Hölder smooth functions satisfying (5), while the hybrid functions satisfying (6) can be similarly analyzed using results from Subsection 3.2.
From practical observations [33], the proximal bundle method works well when the number of inner iterations (i.e., those of Algorithm 2 to solve (3)) stays as a constant much larger than 1 (i.e., that of the subgradient method), say $10$. Recall from Theorem 3.5 that inner complexity is $\tilde{\cal O}\left(\eta L_{\alpha}^{\frac{2}{\alpha+1}}\left(\frac{1}{\delta}\right)^{\frac{1-\alpha}{\alpha+1}}+1\right)$. Since we do not know $\alpha$ and $L_{\alpha}$, we cannot choose a constant stepsize $\eta$ so that the number of inner iterations is close to a desired number such as $10$. Hence, an adaptive stepsize rule is indeed needed.
@@
0:  Let $y_{0}\in\mathbb{R}^{d}$, $\eta_{0}>0$, $\beta_{0}\in(0,1]$, and $\varepsilon>0$ be given.
  for $k=1,2,\cdots$ do
     Call Algorithm 2 with $(y,\eta,\delta)=(y_{k-1},\eta_{k-1},\varepsilon/2)$ and output $(y_{k},\tilde{y}_{k})=(x_{J},\tilde{x}_{J})$.
     if (19) is always true in the execution of Algorithm 2then
        set $\eta_{k}=\eta_{k-1}$;
     else
        set $\eta_{k}=\eta_{k-1}/2$.
     end if
  end for
@@
Assume $f$ is convex and $L_{\alpha}$-Hölder smooth. The following statements hold for APBM:
@@
Proof a) It follows from Lemma 3.1(d) that for every $u\in\mathbb{R}^{d}$
@@
d) This statement immediately follows from (b) and step 3 of Algorithm 3. ∎
@@
Assume $f$ is convex and $L_{\alpha}$-Hölder smooth. If $\eta_{0}\leq\|y_{0}-x_{*}\|^{2}/\varepsilon$, then the iteration-complexity to obtain an $\varepsilon$-solution to (1) is given by
@@
Proof Noting from Lemma 4.1 (d) that $\eta_{k}$ is bounded from below, we know there exists some $\tilde{\eta}\in[\underline{\eta},\eta_{0}]$ such that for some $k_{0}\geq 1$, $\eta_{k}\equiv\tilde{\eta}$ for $k\geq k_{0}$. Thus, it follows from the assumption that $\eta_{0}\leq\|y_{0}-x_{*}\|^{2}/\varepsilon$ that
@@
$2\sum_{k=1}^{n}\eta_{k-1}\left(\min_{1\leq k\leq n}f(\tilde{y}_{k})-f(u)\right)$ $\leq 2\sum_{k=1}^{n}\eta_{k-1}[f(\tilde{y}_{k})-f(u)]$
$\leq\|y_{0}-u\|^{2}-\|y_{n}-u\|^{2}+\varepsilon\sum_{k=1}^{n}\eta_{k-1}.$
@@
for cycles $k\leq k_{0}-1$. Finally, the total iteration-complexity (23) clearly follows from the above two bounds. ∎
Discussion on other universal methods Several universal methods based on the backtracking line-search procedure have been studied in the literature. Paper [48] considers the same Hölder smooth problem (with an additional hybrid function $h$) as in this paper. To finds an $\varepsilon$-solution of (1), the universal primal gradient method proposed in [48] starts from an initial pair $(\hat{x}_{0},\eta_{0})$ and in the $(j+1)$-th iteration searches for a pair $(x_{\eta},\eta)$ satisfying a condition
@@
If the condition (30) is not satisfied, then the method rejects the pair, sets $\eta\leftarrow\eta/2$, and updates $x_{\eta}$ as in (31) with the new $\eta$, otherwise, it accepts the pair and sets $(\hat{x}_{j+1},\eta_{j+1})=(x_{\eta},\eta)$. Two other universal methods are developed in [48], namely, the universal dual gradient method and the universal fast gradient method. Following [48], paper [19] extends the universal fast gradient method to the case of hybrid functions (6). Motivated by the bundle-level method of [30], paper [28] proposes two accelerated variants, i.e., the accelerated bundle-level method and the accelerated prox-level method. The parallel bundle method of [7] is also shown to be universal at the price of running multiple threads.
Paper [34] proposes an adaptive composite subgradient (A-CS) method for solving (1) where $f$ satisfies
@@
for any $\theta>0$. Hence, the Hölder smooth functions (5) considered in this paper are included in the class of functions satisfying (32). More interestingly, a careful look at A-CS of [34] and the universal primal gradient method of [48] reveals that the two methods are identical.
@@
The well-known Nesterov’s accelerated gradient method has been shown in [47] to match the above complexity bound and hence is an optimal method. The accelerated bundle-level method of [28], the universal fast gradient method of [48], and a follow-up work [19] all establish optimal complexity bounds.
On the other hand, the dominant term of bound (23) is its first term and it is only optimal when $\alpha=0$, i.e., $f$ is $L_{0}$-Lipschitz continuous. Motivated by [47], it is possible to develop optimal universal methods based on the accelerated gradient method. This requires accelerated schemes in both PPF and Algorithm 2. Paper [42] proposes an accelerated variant of PPF, which is extended by [3, 20, 14] to obtain optimal $p$-th order methods with convergence rate ${\cal O}(k^{-(3p+1)/2})$ for $p\geq 2$.
We finally note that this paper does not aim to develop the optimal complexity of universal methods; rather, it presents an interesting application of our analysis of Algorithm 2 in the context of universal methods.